OpenStudy (anonymous):

graph each system of equations and state its solution. x+y=6 3x+3y=3

OpenStudy (anonymous):

X+Y=6 =>Y=6-X 3X+3(X-6)=3 =>X+X-6=1 =>2X=7 =>X=7/2 =>Y=6-7/2 =5/2

OpenStudy (anonymous):

there is no solution.

OpenStudy (anonymous):

dhasni, your answer is incorrect.

OpenStudy (anonymous):

would both x's and both y's be the same number?

OpenStudy (anonymous):

did you try graphing the two equations?

OpenStudy (precal):

No solution, because lines are parallel. They have the same slope. m=-1